OpenStudy (lowkey.s):

The elevation of city A is 72 feet above sea level; the elevation of city C is 24 feet above sea level. The elevation of city D is 1 over 4 of the elevation of city C. Part A: Write an expression to represent the elevation, in feet, of city D. Part B: Simplify the expression and explain why the answer is negative or positive.

OpenStudy (lowkey.s):

@Awolflover1 @Addicted2Anime

OpenStudy (lowkey.s):

@kidrah69

OpenStudy (lowkey.s):

@AloneS @ageta

OpenStudy (lowkey.s):

@anglina4196

OpenStudy (lowkey.s):

@Devonhoward15 @Foxythepirate0984 @Frozen15 @pooja195 @PandaNinja101 @Owlcoffee @Ellie201 @LunyMoony @Living4Jesus @YumYum247 @Yinshy @callielovesyhuu @oh_HITZ2523 @gnech @Goldendragon01

OpenStudy (lowkey.s):

I SWEAR I TAG FREAKIN SO MANY PEOPLE AND NO ONE HELP Im basically deaf

OpenStudy (lowkey.s):

@ageta

OpenStudy (lowkey.s):

@sweetburger @study.geek.123 @Shy_Boy @study00 @study00 @study.geek.123 @starryknight

OpenStudy (ageta):

Part A : D = (1/4) * C D = 1/4 * 24 D = 24/4 = 6 ft above sea level Part B : Your answer is positive because D is still above sea level...if it would have been below sea level, then it would have been negative

OpenStudy (anglina4196):

Sorry i have to leave somewhere and i cant answer